# Towards a Scalable Reference-Free Evaluation of Generative Models

### v1

@@ 1 Introduction
On the other hand, a reference-based assessment of generative models may not always be feasible, because the selection of a reference distribution may be challenging in a general learning scenario. For example, in prompt-based generative models where the data are created in response to a user’s input text prompts, the generated data could follow an a priori unknown distribution depending on the specific distribution of the user’s input prompts. Figure 1 illustrates text prompt generated colorful elephant images. Each elephant image corresponds to an unusual color choice, such as luminous yellow or neon pink. Even though reference dataset, such as ImageNet [6] contains elephants, additional elements such as neon colors are not attributed to those images. A proper reference-based evaluation of every user’s generated data would require a distinct reference dataset, which may not be available to the user during the assessment time. Moreover, finding a comprehensive text or video dataset to choose as the reference set would be more difficult compared to image data, because the higher length of text and video samples could significantly contribute to their variety, requiring an inefficiently large reference set to cover all text or video sample types.
Figure 1: Reference-free diversity evaluation of "colorful elephant" samples generated by text-to-image StableDiffusion [13]. FKEA’s top-4 sample clusters and approximated VENDI and RKE are shown.
@@
The discussed challenging scenarios of conducting a reference-based evaluation highlight the need for reference-free assessment methods that remain functional in the absence of a reference dataset. Recently, entropy-based diversity evaluation scores, the $\mathrm{VENDI}$ metric family [1, 14] and $\mathrm{RKE}$ score [2], have been proposed to address the need for reference-free assessment metrics. These scores calculate the entropy of the eigenvalues of a kernel similarity matrix for the generated data. Based on the theoretical results in [2], the evaluation process of these scores can be interpreted as an unsupervised identification of the generative model’s produced sample clusters, followed by the entropy calculation for the frequencies of the detected clusters.
@@
While the $\mathrm{VENDI}$ and $\mathrm{RKE}$ entropy scores provide reference-free assessments of generative models, estimating these scores from generated data could incur significant computational costs. In this work, we show that computing the precise $\mathrm{RKE}$ and $\mathrm{VENDI}$ scores would require at least $\Omega(n^{2})$ and $\Omega(n^{2.376})$ [^1] This computation complexity is the minimum known achievable cost for multiplying $n\times n$ matrices which we prove to lower-bound the complexity of computing matrix-based entropy scores. computations for a sample size $n$, respectively. While the randomized projection methods in [15, 1] can reduce the computational costs to $O(n^{2})$ for a general $\mathrm{VENDI}_{\alpha}$ score, the quadratic growth would be a barrier to the method’s application to large $n$ values. Although the computational expenses could be reduced by limiting the sample size $n$, an insufficient sample size would lead to significant error in estimating the entropy scores.
@@ 3.2 Shift-Invariant Kernels and Random Fourier Features
A kernel function $k$ is called shift-invariant, if there exists a function $\kappa:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$ for every $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}$. Bochner’s theorem proves that a function $\kappa:\mathbb{R}^{d}\rightarrow\mathbb{R}$ will lead to a shift-invariant kernel similarity score $\kappa(x-\mathbf{x}^{\prime})$ between $\mathbf{x},\mathbf{x}^{\prime}$ if and only if its Fourier transform $\widehat{\kappa}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is non-negative everywhere (i.e, $\widehat{\kappa}(\bm{\omega})\geq 0$ for every $\bm{\omega}$). Note that the Fourier transform $\widehat{\kappa}$ is defined as
$\widehat{\kappa}(\bm{\omega})\,:=\,\frac{1}{(2\pi)^{d}}\int{\kappa}(\mathbf{x})\exp\bigl{(}-i\bm{\omega}^{\top}\mathbf{x}\bigr{)}\mathrm{d}\mathbf{x}$
Specifically, Bochner’s theorem shows the Fourier transform $\widehat{\kappa}$ of a normalized shift-invariant kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$, where $\kappa(0)=1$, will be a probability density function (PDF). The framework of random Fourier features (RFFs) [16] utilizes independent samples drawn from PDF $\widehat{\kappa}$ to approximate the kernel function. Here, given independent samples $\bm{\omega}_{1},\ldots,\bm{\omega}_{r}\sim\widehat{\kappa}$, we form the following proxy feature map $\widetilde{\phi}_{r}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2r}$
$\widetilde{\phi}_{r}(\mathbf{x})\,=\,\frac{1}{\sqrt{r}}\Bigl{[}\cos(\bm{\omega}_{1}^{\top}\mathbf{x}),\sin(\bm{\omega}_{1}^{\top}\mathbf{x}),\ldots,\cos(\bm{\omega}_{r}^{\top}\mathbf{x}),\sin(\bm{\omega}_{r}^{\top}\mathbf{x})\Bigr{]}.$ (1)
As demonstrated in [16, 42], the $2r$-dimensional proxy map $\widetilde{\phi}_{r}$ can approximate the kernel function as $k(\mathbf{x},\mathbf{x}^{\prime})=\mathbb{E}_{\bm{\omega}\sim\widehat{\kappa}}\Bigl{[}\cos(\bm{\omega}^{\top}\mathbf{x})\cos(\bm{\omega}^{\top}\mathbf{x}^{\prime})+\sin(\bm{\omega}^{\top}\mathbf{x})\sin(\bm{\omega}^{\top}\mathbf{x}^{\prime})\Bigr{]}\approx\widetilde{\phi}_{r}(\mathbf{x})^{\top}\widetilde{\phi}_{r}(\mathbf{x}^{\prime})$.
@@ 4 Computational Complexity of VENDI & RKE Scores
As discussed, computing $\mathrm{RKE}$ and general $\mathrm{VENDI}_{\alpha}$ scores requires computing the order-$\alpha$ entropy of kernel matrix $\frac{1}{n}K$. Using the standard definition of $\alpha$-norm $\|\mathbf{v}\|_{\alpha}=\bigl{(}\sum_{i=1}^{n}|v_{i}|^{\alpha}\bigr{)}^{1/\alpha}$, we observe that the computation of $\mathrm{VENDI}_{\alpha}$ score is equivalent to computing the $\alpha$-norm $\|\bm{\lambda}\|_{\alpha}$ of the $n$-dimensional eigenvalue vector $\bm{\lambda}=[\lambda_{1},\ldots,\lambda_{n}]$ where $\lambda_{1}\geq\cdots\geq\lambda_{n}$ are the sorted eigenvalues of the normalized kernel matrix $\frac{1}{n}K$.
In the following theorem, we prove that except order $\alpha=2$, which is the $\mathrm{RKE}$ score, computing any other $\mathrm{VENDI}_{\alpha}$ score is at least as expensive as computing the product of two $n\times n$ matrices. Therefore, the theorem suggests that the computational complexity of every member of the VENDI family is lower-bounded by $O(n^{2.367})$ which is the least known cost of multiplying $n\times n$ matrices.
@@
If $\mathrm{VENDI}_{\alpha}(K)$ for $\alpha\neq 2$ is computable by a circuit $\mathcal{C}$ of size $s(n)$ over basis $\mathcal{B}$, then $n\times n$ matrices can be multiplied by a circuit $\mathcal{C}$ of size $O(s(n))$ over basis $\nabla\mathcal{B}|\cup\{+,\times\}$.
@@
The smallest known circuits for multiplying $n\times n$ matrices have size $\Theta(n^{\omega})$, where $\omega\approx 2.367$. Despite tremendous research efforts only minor improvements have been obtained in recent years. There is evidence that $\omega$ is bounded away from 2 for certain classes of circuits [43, 44]. In contrast, $\mathcal{S}_{2}$ is computable in quadratic time $\Theta(n^{2})$ in the basis $B=\{\times,+,\log\}$.
The above discussion indicates that except the $\mathrm{RKE}(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})$, i.e. order-2 Renyi entropy, whose computational complexity is quadratically growing with sample size $\Theta(n^{2})$, the other members of the VENDI family $\mathrm{VENDI}_{\alpha}$ would have a super-quadratic complexity on the order of $\mathcal{O}(n^{2.367})$. In practice, the computation of $\mathrm{VENDI}_{\alpha}$ scores is performed by the eigendecomposition of the $n\times n$ kernel matrix that requires $O(n^{3})$ computations for precise computation and $O(n^{2}M)$ computations using a randomized projection onto an $M$-dimensional space [15, 1].
@@ 5 A Scalable Fourier-based Method for Computing Kernel Entropy Scores
Here, we conisider a shift-invariant kernel matrix $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$ where $\kappa(\mathbf{0})=1$ and propose applying the random Fourier features (RFF) framework [16] to perform an efficient approximation of the RKE and VENDI scores. To do this, we utilize the Fourier transform $\widehat{\kappa}$ that, according to Bochner’s theorem, is a valid PDF, and we independently generate $\bm{\omega}_{1},\ldots,\bm{\omega}_{r}\stackrel{{\scriptstyle\text{\rm iid}}}{{\sim}}\widehat{\kappa}$. Note that in the case of the Gaussian kernel $k_{\text{\rm Gaussian}(\sigma^{2})}$, the corresponding PDF will be an isotropic Gaussian $\mathcal{N}(\mathbf{0},\frac{1}{\sigma^{2}}I_{d})$ with zero mean and covariance matrix $\frac{1}{\sigma^{2}}I_{d}$. Then, we consider the RFF proxy feature map $\widetilde{\phi}_{r}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2r}$ as defined in (1) and define the proxy kernel covariance matrix $\widetilde{C}_{X,r}\in\mathbb{R}^{2r\times 2r}$:
@@
Note that in the above, $\widetilde{\lambda}^{\alpha}_{r,i}$ denotes the $i$th eigenvalue of the $2r\times 2r$ matrix $\widehat{C}_{X,r}$. We remark that the computation of both $\mathrm{FKEA}\text{-}\mathrm{RKE}$ and $\mathrm{FKEA}\text{-}\mathrm{VENDI}_{\alpha}$ can be done by a stochastic algorithm which computes the proxy covariance matrix (2) by summing the sample-based $2r\times 2r$ matrix terms, and then computing the resulting matrix’s Frobenius norm for $\mathrm{RKE}$ score or all the $2r$ matrix’s eigenvalues for a general $\mathrm{VENDI}_{\alpha}$ with $\alpha\neq 2$.
@@
Consider a shift-invariant kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$ where $\kappa(\mathbf{0})=1$. Suppose $\bm{\omega}_{1},\ldots,\bm{\omega}_{r}\sim\widehat{\kappa}$ are independently drawn from PDF $\widehat{\kappa}$. Let $\lambda_{1}\geq\ldots\geq\lambda_{n}$ be the sorted eigenvalues of the normalized kernel matrix $\frac{1}{n}K=\frac{1}{n}\bigl{[}k(\mathbf{x}_{i},\mathbf{x}_{j})\bigr{]}_{n\times n}$. Also, consider the eigenvalues of $\widetilde{\lambda}_{1}\geq\ldots\geq\widetilde{\lambda}_{2r}$ of random matrix $\widetilde{C}_{X,r}$ with their corresponding eigenvectors $\widetilde{\mathbf{v}}_{1},\ldots,\widetilde{\mathbf{v}}_{2r}$. Let $\,\widetilde{\lambda}_{j}=0$ for every $j>2r$. Then, for every $\delta>0$, the following holds with probability at least $1-\delta$:

### v2

@@ 1 Introduction
On the other hand, a reference-based assessment of generative models may not always be feasible, because the selection of a reference distribution may be challenging in a general learning scenario. For example, in prompt-based generative models where the data are created in response to a user’s input text prompts, the generated data could follow an a priori unknown distribution depending on the specific distribution of the user’s input prompts. Figure 1 shows one such example where we compare reference-based diversity scores of regular and colored elephant image samples generated by Stable Diffusion XL [13]. While the diversity of the colored images looks significantly higher to the human eye, the evaluated reference-based FID, Recall, and Coverage metrics do not suggest a higher diversity. As this example suggests, a proper reference-based evaluation of every user’s generated data would require a distinct reference dataset, which may not be available to the user during the assessment time. Moreover, finding a comprehensive text or video dataset to choose as the reference set would be more difficult compared to image data, because the higher length of text and video samples could significantly contribute to their variety, requiring an inefficiently large reference set to cover all text or video sample types.
Figure 1: Reference-based vs. reference-free scores on two datasets of Stable Diffusion XL generated elephant images. FID, Recall, and Coverage scores (colored orange) are reference-based, whereas VENDI and RKE scores (colored blue) are reference-free. Inception.V3 is used as the backbone embedding. Reference-based metrics use ’Indian elephant’ samples in ImageNet as reference data.
@@
The discussed challenging scenarios of conducting a reference-based evaluation highlight the need for reference-free assessment methods that remain functional in the absence of a reference dataset. Recently, entropy-based diversity evaluation scores, the $\mathrm{VENDI}$ metric family [1, 14] and $\mathrm{RKE}$ score [2], have been proposed to address the need for reference-free assessment metrics. These scores calculate the entropy of the eigenvalues of a kernel similarity matrix for the generated data. Based on the theoretical results in [2], the evaluation process of these scores can be interpreted as an unsupervised identification of the generative model’s produced sample clusters, followed by the entropy calculation for the frequencies of the detected clusters. In Figure 1, we observe that the reference-free $\mathrm{VENDI}$ and $\mathrm{RKE}$ scores grow when the generated samples are colored, which is due to the increase in the quantity of identified clusters in the colored case.
@@
While the $\mathrm{VENDI}$ and $\mathrm{RKE}$ entropy scores provide reference-free assessments of generative models, estimating these scores from generated data could incur significant computational costs. In this work, we show that computing the precise $\mathrm{RKE}$ and $\mathrm{VENDI}$ scores would require at least $\Omega(n^{2})$ and $\Omega(n^{2.373})$ [^1] This computation complexity is the minimum known achievable cost for multiplying $n\times n$ matrices which we prove to lower-bound the complexity of computing matrix-based entropy scores. computations for a sample size $n$, respectively. While the randomized projection methods in [15, 1] can reduce the computational costs to $O(n^{2})$ for a general $\mathrm{VENDI}_{\alpha}$ score, the quadratic growth would be a barrier to the method’s application to large $n$ values. Although the computational expenses could be reduced by limiting the sample size $n$, an insufficient sample size would lead to significant error in estimating the entropy scores. As an example on the ImageNet dataset, Figure 7 in the Appendix shows the adverse effects of limiting the sample size on the quality of clusters used in the calculation of the $\mathrm{VENDI}$ scores.
@@ 3.2 Shift-Invariant Kernels and Random Fourier Features
A kernel function $k$ is called shift-invariant, if there exists a function $\kappa:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$ for every $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}$. Bochner’s theorem proves that a function $\kappa:\mathbb{R}^{d}\rightarrow\mathbb{R}$ will lead to a shift-invariant kernel similarity score $\kappa(\mathbf{x}-\mathbf{x}^{\prime})$ between $\mathbf{x},\mathbf{x}^{\prime}$ if and only if its Fourier transform $\widehat{\kappa}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is non-negative everywhere (i.e, $\widehat{\kappa}(\boldsymbol{\omega})\geq 0$ for every $\boldsymbol{\omega}$). Note that the Fourier transform $\widehat{\kappa}$ is defined as
$\widehat{\kappa}(\boldsymbol{\omega})\,:=\,\frac{1}{(2\pi)^{d}}\int{\kappa}(\mathbf{x})\exp\bigl{(}-i\boldsymbol{\omega}^{\top}\mathbf{x}\bigr{)}\mathrm{d}\mathbf{x}$
Specifically, Bochner’s theorem shows the Fourier transform $\widehat{\kappa}$ of a normalized shift-invariant kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$, where $\kappa(0)=1$, will be a probability density function (PDF). The framework of random Fourier features (RFFs) [16] utilizes independent samples drawn from PDF $\widehat{\kappa}$ to approximate the kernel function. Here, given independent samples $\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{r}\sim\widehat{\kappa}$, we form the following proxy feature map $\widetilde{\phi}_{r}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2r}$
$\widetilde{\phi}_{r}(\mathbf{x})\,=\,\frac{1}{\sqrt{r}}\Bigl{[}\cos(\boldsymbol{\omega}_{1}^{\top}\mathbf{x}),\sin(\boldsymbol{\omega}_{1}^{\top}\mathbf{x}),\ldots,\cos(\boldsymbol{\omega}_{r}^{\top}\mathbf{x}),\sin(\boldsymbol{\omega}_{r}^{\top}\mathbf{x})\Bigr{]}.$ (1)
As demonstrated in [16, 42], the $2r$-dimensional proxy map $\widetilde{\phi}_{r}$ can approximate the kernel function as $k(\mathbf{x},\mathbf{x}^{\prime})=\mathbb{E}_{\boldsymbol{\omega}\sim\widehat{\kappa}}\Bigl{[}\cos(\boldsymbol{\omega}^{\top}\mathbf{x})\cos(\boldsymbol{\omega}^{\top}\mathbf{x}^{\prime})+\sin(\boldsymbol{\omega}^{\top}\mathbf{x})\sin(\boldsymbol{\omega}^{\top}\mathbf{x}^{\prime})\Bigr{]}\approx\widetilde{\phi}_{r}(\mathbf{x})^{\top}\widetilde{\phi}_{r}(\mathbf{x}^{\prime})$.
@@ 4 Computational Complexity of VENDI & RKE Scores
As discussed, computing $\mathrm{RKE}$ and general $\mathrm{VENDI}_{\alpha}$ scores requires computing the order-$\alpha$ entropy of kernel matrix $\frac{1}{n}K$. Using the standard definition of $\alpha$-norm $\|\mathbf{v}\|_{\alpha}=\bigl{(}\sum_{i=1}^{n}|v_{i}|^{\alpha}\bigr{)}^{1/\alpha}$, we observe that the computation of $\mathrm{VENDI}_{\alpha}$ score is equivalent to computing the $\alpha$-norm $\|\boldsymbol{\lambda}\|_{\alpha}$ of the $n$-dimensional eigenvalue vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{n}]$ where $\lambda_{1}\geq\cdots\geq\lambda_{n}$ are the sorted eigenvalues of the normalized kernel matrix $\frac{1}{n}K$.
In the following theorem, we prove that except order $\alpha=2$, which is the $\mathrm{RKE}$ score, computing any other $\mathrm{VENDI}_{\alpha}$ score is at least as expensive as computing the product of two $n\times n$ matrices. Therefore, the theorem suggests that the computational complexity of every member of the VENDI family is lower-bounded by $\Omega(n^{2.372})$ which is the least known cost of multiplying $n\times n$ matrices.
@@
If $\mathrm{VENDI}_{\alpha}(K)$ for $\alpha\neq 2$ is computable by a circuit $\mathcal{C}$ of size $s(n)$ over basis $\mathcal{B}$, then $n\times n$ matrices can be multiplied by a circuit $\mathcal{C}$ of size $O(s(n))$ over basis {$\mathcal{B}\cup\nabla\mathcal{B}\cup\{+,\times\}$.
@@
The smallest known circuits for multiplying $n\times n$ matrices have size $\Theta(n^{\omega})$, where $\omega\approx 2.372$. Despite tremendous research efforts only minor improvements have been obtained in recent years. There is evidence that $\omega$ is bounded away from 2 for certain classes of circuits [43, 44]. In contrast, $\mathcal{S}_{2}$ is computable in quadratic time $\Theta(n^{2})$ in the basis $B=\{\times,+,\log\}$.
The above discussion indicates that except the $\mathrm{RKE}(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})$, i.e. order-2 Renyi entropy, whose computational complexity is quadratically growing with sample size $\Theta(n^{2})$, the other members of the VENDI family $\mathrm{VENDI}_{\alpha}$ would have a super-quadratic complexity on the order of $\mathcal{O}(n^{2.372})$. In practice, the computation of $\mathrm{VENDI}_{\alpha}$ scores is performed by the eigendecomposition of the $n\times n$ kernel matrix that requires $O(n^{3})$ computations for precise computation and $O(n^{2}M)$ computations using a randomized projection onto an $M$-dimensional space [15, 1].
@@ 5 A Scalable Fourier-based Method for Computing Kernel Entropy Scores
Here, we consider a shift-invariant kernel matrix $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$ where $\kappa(\mathbf{0})=1$ and propose applying the random Fourier features (RFF) framework [16] to perform an efficient approximation of the RKE and VENDI scores. To do this, we utilize the Fourier transform $\widehat{\kappa}$ that, according to Bochner’s theorem, is a valid PDF, and we independently generate $\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{r}\stackrel{{\scriptstyle\text{\rm iid}}}{{\sim}}\widehat{\kappa}$. Note that in the case of the Gaussian kernel $k_{\text{\rm Gaussian}(\sigma^{2})}$, the corresponding PDF will be an isotropic Gaussian $\mathcal{N}(\mathbf{0},\frac{1}{\sigma^{2}}I_{d})$ with zero mean and covariance matrix $\frac{1}{\sigma^{2}}I_{d}$. Then, we consider the RFF proxy feature map $\widetilde{\phi}_{r}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2r}$ as defined in (1) and define the proxy kernel covariance matrix $\widetilde{C}_{X,r}\in\mathbb{R}^{2r\times 2r}$:
@@
Note that in the above, $\widetilde{\lambda}^{\alpha}_{r,i}$ denotes the $i$th eigenvalue of the $2r\times 2r$ matrix $\widehat{C}_{X,r}$. We remark that the computation of both $\mathrm{FKEA}\text{-}\mathrm{RKE}$ and $\mathrm{FKEA}\text{-}\mathrm{VENDI}_{\alpha}$ can be done by a stochastic algorithm which computes the proxy covariance matrix (2) by summing the sample-based $2r\times 2r$ matrix terms, and then computing the resulting matrix’s Frobenius norm for $\mathrm{RKE}$ score or all the $2r$ matrix’s eigenvalues for a general $\mathrm{VENDI}_{\alpha}$ with $\alpha\neq 2$. Algorithm 1 presents the steps of the FKEA method where the computation needed for the proxy kernel covariance matrix is $O(n)$ and grows only linearly with sample size $n$.
Algorithm 1 FKEA Algorithm for Computing $\mathrm{VENDI}$ and $\mathrm{RKE}$ reference-free scores
1:Input: $n$ datapoints $\mathbf{x}=\{x_{1},\ldots,x_{n}\}$, kernel bandwidth $\sigma^{2}$, RFF dimension $r$
2:Draw i.i.d. samples $\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{r}\sim\hat{\kappa}$triangleright For Gaussian Kernel $\hat{\kappa}\sim\mathcal{N}(0,\frac{1}{\sigma^{2}}I_{d})$
3:Initialize the covariance matrix $\widetilde{C}\leftarrow\mathbf{0}$
4:Compute the covariance matrix:
5:for $i=1$ to $n$ do
6:     Compute the RFF feature for $x_{i}$:
$\widetilde{\phi}_{r}(x_{i})=\frac{1}{\sqrt{r}}\begin{bmatrix}\cos(\boldsymbol{\omega}_{1}^{\top}x_{i}),&\sin(\boldsymbol{\omega}_{1}^{\top}x_{i}),&\cdots,&\cos(\boldsymbol{\omega}_{r}^{\top}x_{i}),&\sin(\boldsymbol{\omega}_{r}^{\top}x_{i})\end{bmatrix}^{\top}$
7:     Update $\widetilde{C}$:
$\widetilde{C}\leftarrow\widetilde{C}+\frac{1}{n}\,\widetilde{\phi}_{r}(x_{i})\,\widetilde{\phi}_{r}(x_{i})^{\top}$
8:end for
9:Perform eigendecomposition on the covariance matrix:
$\{\widetilde{\lambda}_{1},\ldots,\widetilde{\lambda}_{2r}\}\leftarrow\text{Eigendecomposition}(\widetilde{C})$
10:Compute VENDI and RKE metrics using the eigenvalues $\widetilde{\lambda}_{1},\ldots,\widetilde{\lambda}_{2r}$
@@
Consider a shift-invariant kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\kappa(\mathbf{x}-\mathbf{x}^{\prime})$ where $\kappa(\mathbf{0})=1$. Suppose $\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{r}\sim\widehat{\kappa}$ are independently drawn from PDF $\widehat{\kappa}$. Let $\lambda_{1}\geq\ldots\geq\lambda_{n}$ be the sorted eigenvalues of the normalized kernel matrix $\frac{1}{n}K=\frac{1}{n}\bigl{[}k(\mathbf{x}_{i},\mathbf{x}_{j})\bigr{]}_{n\times n}$. Also, consider the eigenvalues of $\widetilde{\lambda}_{1}\geq\ldots\geq\widetilde{\lambda}_{2r}$ of random matrix $\widetilde{C}_{X,r}$ with their corresponding eigenvectors $\widetilde{\mathbf{v}}_{1},\ldots,\widetilde{\mathbf{v}}_{2r}$. Let $\,\widetilde{\lambda}_{j}=0$ for every $j>2r$. Then, for every $\delta>0$, the following holds with probability at least $1-\delta$: